The brass rod is 2m long at a certain temperature. What is the length for a temperature rise of 100k, if the expansivity of brass is 18 × 10^-6k^-1?


1
Expert's answer
2021-02-07T19:09:05-0500

By definition, the gain in length will be:


"\Delta L = \alpha L_0\Delta T"

where "\alpha = 18\times 10^{-6}K^{-1}" is the expansivity, "L_0 = 2m" is the initial length, and "\Delta T = 100K" is the change in temperature. Thus, obtain:


"\Delta L = 18\times 10^{-6}\cdot 2\cdot 100 = 3.6\times 10^{-3} = 0.0036m"

Hence, the final length will be:


"L = L_0 + \Delta L = 2m + 0.0036m = 2.0036m"

Answer. 2.0036m.